# Unsupervised Multimodal Graph-based Model for Geo-social Analysis

### v1

@@ 4.1 Experimental Setup
Network Training and Parametrization: The models was trained in an unsupervised manner, without access to ground truth topic or cluster labels, using the compound loss, as detailed in the Section 3.2. Input features consisted of SBERT-generated semantic embeddings for textual content and corresponding geolocation coordinates for graph construction. These modalities were used to define semantic and geographic edges in both the MonoGraph and MultiGraph formulations. Training was performed using the Adam optimizer with an initial learning rate of 1e-3. Models were trained for 50 epochs, with early stopping based on the stability of the coherence score over a held-out validation set. In the MultiGraph setup, we applied a dropout rate of 0.3 and used two Torch GCN layers per branch. The multi-head attention mechanism used for embedding fusion consisted of 4 heads and incorporated layer normalization between fusion stages. To efficiently manage memory, avoid over-sized graphs, and enable consistent training across varying dataset sizes, we adopted a strided chunking protocol. In specific, each dataset was partitioned into multiple non-overlapping chunks using a fixed stride value. Given a stride $s$, each chunk consisted of every $s$-th entry from the original dataset, starting from a different offset. Formally, we generated $s$ subsets $\{\text{chunk}_{i}\}_{i=1}^{s}$ such that:
$\text{chunk}_{i}=\text{Input}_{\text{gdf}}[i::s]$
This approach preserved the overall diversity and temporal structure of the data while enabling the model to process each chunk independently. The protocol ensures that all data points are covered, and total recovery of the original dataset was verified before training. Each chunk was used sequentially as input to the model in a looped training cycle. Hyperparameter tuning was carried out via a random search over a validation subset of the datasets. Ultimately, we set the contrastive, coherence, and alignment loss weights to $\alpha$ = 0.8, $\beta$ = 0.2, and $\gamma$ = 0.1, respectively. These choices are based on a combination of domain knowledge, empirical tuning, and the relative importance of each loss component for the task at hand. For the contrastive loss, we set the temperature parameter $\tau$ = 0.5. All experiments were executed on an NVIDIA H100 GPU.
@@ 4.2 Clustering Capability Evaluation
To evaluate the models’ capacity for generating semantically coherent representations, we conducted clustering assessments on both the original SBERT input embeddings and the output embedded features produced by each model. Following the methodology outlined in Section 3.2, we computed two key metrics: intra-cluster similarity, reflecting semantic cohesion within clusters, and inter-cluster similarity, indicating the degree of separation between them. Ideally, effective clustering is characterized by high intra-cluster cohesion and low inter-cluster similarity. Table 2 summarizes the results across the four disaster-specific datasets. The input embeddings serve as a baseline and consistently exhibit moderate cohesion but relatively poor separation, suggesting that the raw SBERT representations lack strong semantic structure. Both MonoGraph and MultiGraph yield substantial improvements. Across all datasets, inter-cluster similarity drops significantly, while intra-cluster cohesion improves. MultiGraph achieves the highest cohesion values on all datasets and attains the lowest (i.e., most negative) inter-cluster similarity scores for the Chile wildfires and Napa earthquake datasets, suggesting well-isolated and internally consistent clusters. On the Ahr Valley floods and Hurricane Harvey datasets, MonoGraph shows better (lower) inter-cluster similarity values, although its intra-cluster scores are slightly lower than MultiGraph’s, yet still markedly better than the baseline. These trends are further reinforced by the visualizations in Fig. 2. The input similarity matrix (top-left) exhibit less well-defined structure, whereas the output matrices (middle-left and bottom-left) display sharper contrast and more structured groupings. Likewise, the 3D cluster plots (right) reveal a transition from scattered distributions in the input space to more compact and semantically coherent clusters after model transformation. Low inter-cluster similarity values indicate strong dissimilarity between clusters, reflecting effective cluster separation. In summary, both models, especially MultiGraph, demonstrate a strong ability to reorganize the input features space in ways that enhance clustering quality. This improvement is crucial for downstream tasks such as topic modeling, trend detection, and spatial-semantic analysis.
4.3 Benchmark and Baseline Comparison
The objective of this evaluation was to quantitatively assess the models’ performance and their capability to generate high-quality clusters across diverse disaster scenarios. For benchmarking purposes, we compared our proposed models against the multimodal pipeline presented by Hanny et al. [hanny2025], as well as the baseline sequential workflow introduced in this study. To rule out the effect of time and sentiment features in the multimodal pipeline, we set the sentiment and time weights in the pipeline to zero during the evaluation. To ensure fair comparison, we used the same exact datasets as used in [hanny2025]. For embedding generation, we utilized the English-only all-MiniLM-L12-v2 SBERT model for the 2014 Napa earthquake and 2017 Hurricane Harvey datasets. Conversely, for the 2021 Ahr Valley flood and 2023 Chile wildfires datasets, we adopted the multilingual distiluse-base-multilingual-cased-v1 model, consistent with the approach in [hanny2025]. To ensure a congruent comparison across models, we chose Topic Quality (TQ), a widely used metric in the community. TQ effectively evaluates the coherence and relevance of topics, making it a reliable measure for assessing the quality of topic models. It is computed based on the top 10 keywords per topic and defined as the product of Topic Coherence (TC) and Topic Diversity (TD). We employed a TF-IDF-based keyword extraction method: For each cluster, the corresponding subset of documents was aggregated, and TF-IDF scores were calculated over this localized corpus. The top 10 highest-scoring terms were selected as representative keywords.
To compute the TQ score, we first constructed a document-term matrix using CountVectorizer, then estimated word and word-pair frequencies across the full corpus. For each topic, pairwise Positive Pointwise Mutual Information (PPMI) between all keyword pairs was calculated and normalized by the negative log joint probability to obtain the normalized variant (NPPMI) [bouma2009] [jurafsky2023]. Formally, the normalized PPMI between two words $w_{i}$ and $w_{j}$ is given by
$\text{NPPMI}(w_{i},w_{j})=\frac{\max(0,\log\frac{P(w_{i},w_{j})}{P(w_{i})P(w_{j})})}{-\log P(w_{i},w_{j})},$
@@ 4.3 Benchmark and Baseline Comparison
where probabilities are estimated from co-occurrence counts over the corpus. The coherence of a topic is the average NPPMI over all $\binom{10}{2}$ word pairs in the topic, and the overall Topic Coherence is the mean across all valid topics. Topic Diversity quantifies how distinct the discovered topics are by measuring the proportion of unique keywords across all topics. It is calculated as
@@
where $U$ denotes the number of unique words in the top-$k$ keywords across all topics, $k$ is the number of keywords per topic, and $|T|$ is the total number of topics. This penalizes repeated keywords across topics and rewards diversity in word usage. Finally, Topic Quality is computed as the product of the two metrics:
@@
which favors models that produce both semantically coherent and lexically diverse clusters. All metrics are computed using the same preprocessing pipeline for consistency across evaluated methods.
@@ 4.4 TQ Results and Analysis
The Topic Quality (TQ) scores reported in Table 3 clearly demonstrate the effectiveness of our proposed models in generating high-quality topic clusters across diverse disaster-related datasets. The MultiGraph consistently achieves the highest TQ scores in three out of four scenarios, the Ahr Valley floods (0.415), Hurricane Harvey (0.384), and the Napa earthquake (0.195), demonstrating its robustness and adaptability to datasets with varying linguistic, geographic, and topical characteristics. This superior performance can be attributed to its cross-attention mechanism, which enables selective regulation of individual modalities and more effective capture of both global semantic patterns and local topical signals. On the Chile wildfires dataset, however, the MonoGraph slightly outperforms the MultiGraph (0.169 vs. 0.161), with both models performing comparably well overall. This suggests that while the MultiGraph generally excels, the benefits of increased model complexity may vary depending on the nature and quality of the data. In particular, contexts characterized by fewer high-quality labeled signals or more multilingual noise may favor the simpler and more direct MonoGraph architecture, which consistently outperforms baseline methods across all datasets and yields stable, focused clusters despite its relative simplicity.
Table 3: Topic quality scores (TQ) for outputs from our two models (MonoGraph and MultiGraph), the method by Hanny et al. [hanny2025], and the baseline sequential workflow across four datasets.

### v2

@@ 4.3 Model Configuration and Training
All models were trained in an unsupervised manner without access to ground-truth topic or cluster labels, using the compound objective described in Section 3.2. Input features consisted of SBERT-generated semantic embeddings and associated geolocation coordinates, which were used to construct semantic and geographic graphs in both the MonoGraph and MultiGraph formulations. Training was performed using the Adam optimizer with an initial learning rate of $1\times 10^{-3}$. Models were trained for 50 epochs with early stopping based on coherence stability on a held-out validation subset. In the MultiGraph configuration, each branch employed two GCN layers with a dropout rate of 0.3. Fusion was performed via multi-head attention with four heads, and layer normalization was applied between fusion stages to stabilize training.
4.4 Scalability via Strided Chunking
To scale graph construction to large corpora, we adopt strided chunking during preprocessing. Given a stride $s$, the dataset is partitioned into $s$ partially overlapping subsets defined as
$\text{chunk}_{i}=\text{Input}_{\mathrm{gdf}}[i::s],\quad i=1,\dots,s.$
This construction ensures that each node appears in exactly one position per stride cycle while nodes near chunk boundaries are shared across multiple chunks, preserving local semantic and geographic neighborhoods. Chunk overlap prevents artificial fragmentation of neighborhoods that could otherwise arise from independent graph construction. As a result, node embeddings remain consistent across chunks, and striding affects computational efficiency only, not the learned representation or clustering behavior. Full data coverage was verified prior to training, and all chunks were processed sequentially under identical optimization settings. Empirically, varying the stride within reasonable bounds did not result in meaningful changes in the embedding geometry or TQ, confirming that strided chunking functions strictly as a scalability mechanism and does not bias embedding geometry or topic structure. Stride values are selected based on memory constraints rather than dataset-specific tuning, and do not affect neighborhood structure or learned representations.
4.5 Baseline Configuration
The primary baseline for all experiments is the sequential semantic workflow introduced in this study. It operates exclusively on semantic embeddings and does not incorporate graph-based modeling, geographic information, multimodal fusion, or contrastive learning. This configuration serves as a minimal reference point and enables direct attribution of performance differences to the proposed graph construction, semantic–geographic integration, and contrastive objectives. In addition, we include the JSTTS multimodal pipeline proposed by Hanny et al. [hanny2025] as an external reference method. Although temporal and sentiment features are part of the original JSTTS formulation, these modalities are intentionally excluded from all primary experiments in this work. This design choice applies uniformly across all evaluated models and is motivated by the goal of isolating the methodological contribution of semantic–geographic graph modeling and contrastive learning. Incorporating auxiliary temporal or affective signals would alter the input space and introduce additional information that is orthogonal to the methodological focus of this study, thereby confounding attribution of performance differences. Nonetheless, for completeness and transparency, the behavior of JSTTS under its original full-modality configuration is reported separately in the evaluation section. These results are provided for contextual reference only and are not used for direct comparison with the proposed models.
To ensure comparability with prior work, we use the same datasets and embedding backbones as in [hanny2025] across all experiments, including clustering and topic quality evaluation. Specifically, the English all-MiniLM-L12-v2 model is used for the Napa earthquake and Hurricane Harvey datasets, while the multilingual distiluse-base-multilingual-cased-v1 model is used for the Ahr Valley floods and Chile wildfires datasets. All models are evaluated using identical preprocessing pipelines, clustering procedures, and keyword extraction methods.
5 Evaluation
This section evaluates the proposed models from two complementary perspectives. First, we assess their ability to produce semantically coherent and well-separated clusters in the embedding space. Second, we benchmark topic modeling performance using TQ against a sequential baseline and an established multimodal pipeline. Together, these evaluations quantify both representation quality and downstream topic modeling effectiveness across diverse disaster scenarios.
@@ 5.1 Clustering Capability Evaluation
To evaluate the models’ capacity for generating semantically coherent representations, we conducted clustering assessments on both the original input embeddings and the output embedded features produced by each model. Following the methodology outlined in Section 3.2, we computed two complementary metrics: intra-cluster similarity, reflecting semantic cohesion within clusters, and inter-cluster similarity, indicating the degree of separation between clusters. Effective clustering is characterized by high intra-cluster cohesion and low inter-cluster similarity.
Table 2 summarizes the results across four disaster-specific datasets. The raw input embeddings serve as a reference baseline and consistently exhibit moderate intra-cluster cohesion but relatively poor inter-cluster separation, indicating that semantic structure alone is insufficient for robust clustering. Both MonoGraph and MultiGraph substantially improve clustering behavior across all datasets. Inter-cluster similarity decreases markedly, while intra-cluster cohesion increases, indicating tighter and better-separated clusters. MultiGraph achieves the highest intra-cluster cohesion across all datasets and yields the lowest (i.e., most negative) inter-cluster similarity for the Chile wildfires and Napa earthquake datasets, suggesting particularly well-isolated and internally consistent clusters. For the Ahr Valley floods and Hurricane Harvey datasets, MonoGraph attains slightly lower inter-cluster similarity values, although its intra-cluster cohesion remains marginally below that of MultiGraph while still substantially outperforming the baseline.
These quantitative findings are reinforced by the visualizations in Fig. 2. The input similarity matrix exhibits diffuse structure, whereas the learned representations display sharper contrast and clearer block patterns. Likewise, the 3D visualizations reveal a transition from scattered point clouds to compact and well- separated clusters. Overall, both models, and especially MultiGraph, demonstrate a strong ability to reorganize the embedding space in a way that enhances clustering quality, which is critical for downstream topic modeling and spatial-semantic analysis.
5.2 Topic Quality Benchmarking
This section evaluates the semantic quality of the topics discovered by the proposed models under controlled and comparable conditions. We first describe the evaluation protocol and baseline configurations used for TQ assessment, then formally define the TQ metric, and finally present quantitative results across all datasets.
5.2.1 Evaluation Protocol
Following the baseline definitions established in Section 4.5, TQ evaluation is conducted under controlled and directly comparable conditions. All models are evaluated using semantic embeddings as the common input modality. Both the multimodal models additionally incorporate geographic information, while the BERTopic baseline operates solely on semantic representations, reflecting its original formulation. In addition to the primary benchmarking experiments, we report TQ scores for the JSTTS pipeline proposed by Hanny et al. [hanny2025] in its original intended configuration, with all available modalities, including semantic, spatial, temporal, and sentiment features, enabled (reported in Table 3). As already mentioned in Section 4.5, introducing auxiliary information that is orthogonal to the methodological focus of this study. As a result, these results are not treated as directly comparable benchmarks, but are reported for transparency and contextual reference, illustrating the relative scale of TQ achieved by modular multimodal pipelines.
5.2.2 Topic Quality Metric
Topic quality is a widely adopted metric for evaluating topic models, as it jointly captures semantic coherence and lexical diversity. It is defined as the product of Topic Coherence (TC) and Topic Diversity (TD), computed using the top-$k$ keywords per topic (with $k_{\text{means}}=15$ in all experiments). To extract representative keywords, documents assigned to each cluster are aggregated and a TF–IDF vectorizer is applied locally. The top-$k$ terms with the highest TF–IDF scores are selected as topic descriptors. Topic Coherence is computed using normalized Positive Pointwise Mutual Information (NPPMI) [bouma2009, jurafsky2023]. Given two words $w_{i}$ and $w_{j}$, NPPMI is defined as:
$\text{NPPMI}(w_{i},w_{j})=\frac{\max\left(0,\log\frac{P(w_{i},w_{j})}{P(w_{i})P(w_{j})}\right)}{-\log P(w_{i},w_{j})},$
@@ 5.2.2 Topic Quality Metric
where probabilities are estimated from word co-occurrence counts over the full corpus. The coherence of a topic is computed as the average NPPMI over all $\binom{k}{2}$ keyword pairs, and the overall Topic Coherence is obtained by averaging across all valid topics.
Topic Diversity measures how distinct the discovered topics are by quantifying the proportion of unique keywords across all topics:
@@
where $U$ denotes the number of unique words among the top-$k$ keywords across all topics, and $|T|$ is the total number of topics. This formulation penalizes repeated keywords and rewards lexical diversity. Finally, TQ is defined as:
@@
favoring models that produce topics that are both semantically coherent and lexically diverse. All components of TQ are computed using the same preprocessing and keyword extraction pipeline for all evaluated methods.
Table 3: Topic Quality(TQ) achieved by the JSTTS pipeline when all modalities (semantic, spatial, temporal, and sentiment) are enabled, as originally specified.
       Dataset        JSTTS TQ
       Ahr Valley Floods        0.165
       Hurricane Harvey        0.191
       Chile Wildfires        0.142
       Napa Earthquake        0.081
@@ 5.3 TQ Results and Analysis
The TQ scores reported in Table 4 clearly demonstrate the effectiveness of our proposed models in generating high-quality topic clusters across diverse disaster-related datasets. The MultiGraph consistently achieves the highest TQ scores in three out of four scenarios, the Ahr Valley floods (0.415), Hurricane Harvey (0.384), and the Napa earthquake (0.195), demonstrating its robustness and adaptability to datasets with varying linguistic, geographic, and topical characteristics. This superior performance can be attributed to its cross-attention mechanism, which enables selective regulation of individual modalities and more effective capture of both global semantic patterns and local topical signals. On the Chile wildfires dataset, however, the MonoGraph slightly outperforms the MultiGraph (0.169 vs. 0.161), with both models performing comparably well overall. This suggests that while the MultiGraph generally excels, the benefits of increased model complexity may vary depending on the nature and quality of the data. In particular, contexts characterized by fewer high-quality labeled signals or more multilingual noise may favor the simpler and more direct MonoGraph architecture, which consistently outperforms baseline methods across all datasets and yields stable, focused clusters despite its relative simplicity.
Table 4: TQ for outputs from our two models (MonoGraph and MultiGraph), the method by Hanny et al. [hanny2025], and the baseline sequential workflow across four datasets.
@@
6 Ablation, Sensitivity, and Stability Analysis
To provide a systematic understanding of the proposed multimodal graph-based topic modeling framework, we conduct a comprehensive set of ablation, sensitivity, and stability experiments. These analyses are designed to: (i) isolate the contribution of individual loss components and architectural choices, and (ii) assess the robustness of the model with respect to hyperparameter settings and stochastic initialization. All experiments are conducted on the Hurricane Harvey dataset using both the MonoGraph and MultiGraph variants. Performance is evaluated using Topic Quality, which jointly reflects topic coherence and diversity. Unless otherwise stated, results reported in this section are averaged over the corresponding experimental settings, and summary statistics are provided in the form of Mean, standard deviation (STD), and 95% Confidence Interval (CI) where applicable.
Table 5: Loss ablation results based on TQ. ✓ indicates the loss is included, while $\times$ indicates it is excluded.
   Model    $\mathcal{L}_{c}$    $\mathcal{L}_{h}$    $\mathcal{L}_{a}$    Removed Loss    TQ
   MonoGraph    ✓    ✓    ✓    None (Full)    0.368
   MonoGraph    $\times$    ✓    ✓    Contrastive    0.205
   MonoGraph    ✓    $\times$    ✓    Coherence    0.357
   MonoGraph    ✓    ✓    $\times$    Alignment    0.353
   MultiGraph    ✓    ✓    ✓    None (Full)    0.384
   MultiGraph    $\times$    ✓    ✓    Contrastive    0.218
   MultiGraph    ✓    $\times$    ✓    Coherence    0.362
   MultiGraph    ✓    ✓    $\times$    Alignment    0.366
6.1 Loss Ablation Analysis
We analyze the contribution of each loss component, contrastive loss ($\mathcal{L}_{c}$), coherence loss ($\mathcal{L}_{h}$), and alignment loss ($\mathcal{L}_{a}$), by removing one term at a time while keeping all other hyperparameters fixed at their baseline configuration. Results are summarized in Table 5. Removing the contrastive loss leads to a substantial degradation in performance for both model variants, with TQ dropping sharply relative to the full configuration. This indicates that contrastive learning provides the primary discriminative signal necessary for effective topic separation. In contrast, removing either the coherence or alignment loss results in only moderate reductions in TQ, suggesting that these components act as stabilizing regularizers rather than dominant drivers of clustering quality. Overall, the ablation results confirm that the loss formulation is well balanced: contrastive learning establishes topic structure, while coherence and alignment losses refine and stabilize the learned representations.
Table 6: Sensitivity analysis of loss weights based on Topic TQ.
     Setting      MonoGraph      MultiGraph
     $\alpha=0.2$      0.308      0.346
     $\alpha=0.5$      0.351      0.355
     $\alpha=0.8$      0.368      0.384
     $\beta=0.1$      0.344      0.363
     $\beta=0.2$      0.368      0.384
     $\beta=0.5$      0.365      0.354
     $\lambda=0.05$      0.361      0.367
     $\lambda=0.1$      0.368      0.384
     $\lambda=0.2$      0.349      0.362
     Mean      0.354      0.367
     STD      0.019      0.014
     95% CI      $\pm$0.013      $\pm$0.009
Table 7: Sensitivity analysis for contrastive temperature $\tau$ and number of clusters ($k_{\text{means}}=15$) using TQ.
   Parameter    Value    MonoGraph TQ    MultiGraph TQ
   $\tau$    0.2    0.314    0.359
   $\tau$    0.5    0.368    0.384
   $\tau$    1.0    0.315    0.359
   $k_{\text{means}}=15$    10    0.341    0.351
   $k_{\text{means}}=15$    15    0.368    0.384
   $k_{\text{means}}=15$    20    0.364    0.370
   Mean    –    0.345    0.368
   STD    –    0.026    0.017
   95% CI    –    $\pm$0.021    $\pm$0.014
Table 8: Stability across random seeds computed over independent runs.
  Model   Mean TQ   STD   95% CI   Seeds
  MonoGraph   0.3667   0.0025   $\pm$0.0062   {42,1,2,3}
  MultiGraph   0.3625   0.0034   $\pm$0.0084   {42,1,2,3}
6.2 Hyperparameter Sensitivity Analysis
We assess robustness to hyperparameter selection by independently varying key parameters while keeping all others fixed at their baseline configuration. Table 6 summarizes TQ under different loss-weight configurations. For MonoGraph, the mean TQ across all tested settings is $0.354$ with a standard deviation of $0.019$, while MultiGraph achieves a mean of $0.367$ with a standard deviation of $0.014$. The relatively small dispersion indicates that performance is stable under moderate variations of loss weights, with the baseline configuration consistently yielding the highest TQ. Sensitivity with respect to the contrastive temperature $\tau$ and the number of clusters ($k_{\text{means}}=15$) is reported in Table 7. TQ exhibits a unimodal trend with respect to $\tau$, where intermediate values provide optimal separation of positive and negative pairs. Across all tested values, the standard deviation remains below $0.03$ for both models, indicating limited sensitivity. Similarly, performance remains stable across different values of ($k_{\text{means}}$), with peak TQ achieved at the baseline configuration $k_{\text{means}}=15$. It is also worth noting that while the absolute topic granularity varies with the choice of $k_{\text{means}}$, the relative performance ranking between MonoGraph and MultiGraph remains stable across the tested range. These results demonstrate that the proposed framework is robust to reasonable variations in both optimization and clustering hyperparameters, and does not rely on finely tuned parameter choices.
Table 9: Comparison of fusion strategies for multimodal integration based on TQ.
     Model      Fusion Strategy      TQ
     MultiGraph      Concatenation      0.225
     MultiGraph      Attention      0.384
     MultiGraph      Concatenation + MLP      0.165
6.3 Random Seeds Stability and Fusion Strategy
To evaluate robustness to stochastic initialization, we repeat the full baseline configuration using multiple random seeds. Table 8 reports the mean TQ, standard deviation, and 95% confidence interval. Both models exhibit low variance across runs, indicating that performance is stable and reproducible. We further compare attention-based fusion with alternative fusion strategies for the MultiGraph model, including plain concatenation and concatenation augmented with a multi-layer perceptron (MLP). As shown in Table 9, attention-based fusion yields substantially higher TQ than both concatenation-based variants, indicating that adaptive, context-aware weighting is more effective for integrating heterogeneous modalities than static or prior-driven fusion schemes.
6.4 Architectural and graph construction choices
In addition to the reported ablation and sensitivity analyses, we conducted preliminary validation experiments to assess the impact of architectural depth and graph construction choices on model behavior. In particular, we varied the number of GCN layers and attention heads within a small range (1–3) and observed no consistent or systematic changes in TQ across datasets. Based on this observation, these architectural parameters are fixed in all reported experiments in order to reduce the hyperparameter search space and improve reproducibility. We also explored alternative graph construction strategies, including global similarity and distance-based thresholding. These approaches were found to be sensitive to dataset scale and density, often requiring dataset-specific tuning to maintain stable sparsity levels. To avoid this dependency, we adopt a consistent top-$k$ neighborhood strategy for both semantic and geographic graphs throughout the paper. This choice yields predictable graph sparsity across datasets and allows performance differences to more directly reflect the proposed multimodal fusion mechanism and learning objectives rather than incidental graph construction effects.
6.5 Overall Model Behavior
Overall, the ablation, sensitivity, and stability analyses demonstrate that the proposed framework is robust across a wide range of configurations. Performance consistently peaks at the baseline configuration ($\alpha=0.8$, $\beta=0.2$, $\lambda=0.1$, $seed=42$, $k_{\text{means}}=15$, and $\tau=0.5$) while maintaining low variance across random seeds. These results indicate that the reported improvements are not artifacts of hyperparameter tuning or stochastic effects, but reflect stable and reproducible behavior of the proposed model.